A 105.2 mL sample of 1.00 M NaOH is mixed with 52.6 mL of 1.00 M H2SO4 in a large Styrofoam coffee cup; the cup is fitted with a lid through which passes a calibrated thermometer. The temperature of each solution before mixing is 23.05 °C. After adding the NaOH solution to the coffee cup and stirring the mixed solutions with the thermometer, the maximum temperature measured is 30.70 °C. Assume that the density of the mixed solutions is 1.00 g/mL, that the specific heat of the mixed solutions is 4.18 J/(g·°C), and that no heat is lost to the surroundings.

Calculate the enthalpy change per mole of H2SO4 in the reaction.
Answer #1

no of moles of NaOH = molarity*volume in L

                                = 1*0.1052   = 0.1052moles

no of moles of H2SO4 = molarity*volume in L

                                   = 1*0.0526   = 0.0526moles

H2SO4(aq) + 2NaOH(aq) --------> Na2SO4(aq) + 2H2O(l)

total volume of solution = 105.2 + 52.6   = 157.8ml

mass of solution = volume * density

                       = 157.8 *1 = 157.8g

q = mc\DeltaT

    = 157.8*4.18*(30.70-23.05)

     = 5046J

\DeltaH = -q = -5045J

\DeltaH    = -5045J/0.0526mole

            = -95913J/mole

           = -95.913KJ/mole >>>>>>answer